OpenStudy (codyhelton):

The numbers 77 and 791 are relatively prime. True False

OpenStudy (codyhelton):

@Mariamikayla

OpenStudy (mariamikayla):

77=7*11 791= 7*113 It's false because they have in common number 7 as a divisor.

OpenStudy (mahoganie.carson):

Maria right it is flase Cody.